# FSPGD: Rethinking Black-box Attacks on Semantic Segmentation

### v1

@@ B Extended Experimental Results
FSPGD(Ours) 3.28 11.44 13.47 21.67
@@ C Additional Examples for Qualitative Evaluation
Figure 5: Visualization of clean image, attacked images, and output predictions on Pascal VOC 2012. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 (second row), PSPNet-Res101 (third row), and FCNVGG16 (fourth row) are used as target models. (a) Clean image, (b) PGD [34], (c) SegPGD [16], (d) CosPGD [1], (e) FSPGD (Ours).
Figure 6: Visualization of clean image, attacked images, and output predictions on Pascal VOC 2012. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 (second row), PSPNet-Res101 (third row), and FCNVGG16 (fourth row) are used as target models. (a) Clean image, (b) PGD [34], (c) SegPGD [16], (d) CosPGD [1], (e) FSPGD (Ours).
Figure 7: Visualization of clean image, attacked images, and output predictions on Cityscapes. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 (second row), PSPNet-Res101 (third row), and FCNVGG16 (fourth row) are used as target models. (a) Clean image, (b) PGD [34], (c) SegPGD [16], (d) CosPGD [1], (e) FSPGD (Ours).
Figure 8: Visualization of clean image, attacked images, and output predictions on Cityscapes. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 (second row), PSPNet-Res101 (third row), and FCNVGG16 (fourth row) are used as target models. (a) Clean image, (b) PGD [34], (c) SegPGD [16], (d) CosPGD [1], (e) FSPGD (Ours).
@@ D Detailed Experimental Results for Ablation Studies
This section presents the quantified experimental results used to make the graphs (i.e. Figs. 4,  5, and 6 in main paper) in ablation studies discussed in Sec. 4.3 and provides a more detailed explanation of these results. Additionally, it elaborates on ablation study findings that were not included in the main text due to space constraints.
@@ D.1 Performance comparison based on $\tau$ value
The proposed method includes a user-defined parameter, $\tau$, which is used to build the mask $\textrm{M}_{B}$. Since the $\tau$ value affects the attack performance, we conducted extensive experiments to compare the results. As shown in Table 3, the attack performance varies slightly depending on the $\tau$ value. Notably, although performance fluctuates with different $\tau$ values, it consistently outperforms conventional techniques shown in Tables 1 and 2 in main paper. We calculated the average performance for each $\tau$ value and selected cos($\pi/3$) as the optimal value, as it achieved the highest average performance.
@@
$\pi/2$ 3.58 26.48 29.02 28.80
@@ D.2 Performance comparison based on $\lambda$ value
The proposed loss function consists of two loss terms, $L_{ex}$ and $L_{in}$. To examine the effect of each loss term, we build the performance graph as depicted in Fig. 5 in main paper. Here, we provide a detailed numerical explanation of the experimental results used to generate Fig. 5 in main paper, along with additional results for new loss term combinations not included in Fig. 5. Table 4 summarizes the experimental results on the Pascal VOC 2012 dataset. As shown in the Table 4, the performance of the proposed method varies depending on how the two loss terms are combined. As discussed in the main text, simply adding the two loss terms can result in a compromise, leading to lower performance compared to using $L_{ex}$ alone. To investigate this performance degradation, we conducted experiments with different ratios, such as $L_{ex}+0.5L_{in}$ and $L_{ex}+0.1L_{in}$. The results, as summarized in the Table 4, show that performance varies depending on the source model; for instance, when PSPNet-Res50 is the source model, performance was lower compared to using $L_{ex}$ alone, but when DeepLabv3-Res50 was used, performance improved. To address this issue of performance variation across source models, we proposed a dynamic $\lambda_{t}$ that adjusts with t, and this method demonstrated the best performance overall.
@@
  Source Models $\lambda$ Target Models
Source Model PSPRes101 DVRes101 FCNVGG16
@@
$\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$ 3.44 22.02 16.57 19.36
@@ D.3 Performance comparison based on layer location
Unlike conventional methods, the proposed method performs attacks by leveraging intermediate-layer features, making it the first approach to introduce intermediate-layer attacks in the field of semantic segmentation. As such, unlike intermediate-layer attack methods in image classification, there is no prior research on which layer is optimal for attacks in semantic segmentation. To address this, we conducted extensive experiments by attacking various layers of the encoder and summarized the results. Tables 5 and 6 present the intermediate-layer attack performance for encoders ResNet50 and ResNet101, respectively. Attacking the later layers of the encoder (i.e., layer 4_2) results in strong performance on the source model but poor performance on the target models. In contrast, attacking the middle layers demonstrates reasonable attack performance on the source model while also achieving high transferability. Therefore, as the proposed method aims to enhance transferability, we chose to attack the middle layers.

### v2

@@ B Extended Experimental Results
FSPGD(Ours) 3.28 11.42 13.45 21.49
Table 2: Attack performance comparison on Cityscapes in terms of mIoU. Lower mIoU means better performance and bold numbers denote the best mIoU values for each experimental setup
[…]
Source Models Attack Method Source PSP DV3 Segformer Maskformer
Model Res101 Res101 MiT-B0 Swin-S
Clean Images 64.62 / 65.90 65.65 67.16 60.58 68.24
PSP PGD [36] 1.83 18.80 19.35 48.92 59.66
Res50 SegPGD [17] 1.38 18.26 19.34 49.83 60.41
CosPGD [1] 0.07 24.90 26.65 50.31 60.53
DAG [53] 23.52 36.76 33.47 50.24 60.64
NI [32] 1.62 15.07 17.07 44.43 50.09
DI [54] 1.92 17.60 21.57 52.12 54.81
TI [11] 1.64 28.39 34.07 51.91 58.70
FSPGD (Ours) 0.93 5.12 3.29 41.30 47.30
DV3 PGD [36] 2.00 22.19 22.06 50.28 60.64
Res50 SegPGD [17] 0.96 22.20 22.51 50.59 60.24
CosPGD [1] 0.01 25.43 27.22 50.48 59.86
DAG [53] 36.54 39.01 35.98 51.59 60.19
NI [32] 1.55 16.65 18.26 45.76 49.89
DI [54] 2.32 19.87 23.61 52.63 55.32
TI [11] 1.48 31.93 35.45 52.77 59.56
FSPGD (Ours) 1.27 6.09 3.78 40.74 47.30
Source Models Attack Method Source PSP DV3 Segformer Maskformer
Model Res50 Res50 MiT-B0 Swin-S
Clean Images 65.65 / 67.16 64.62 65.90 60.58 68.24
PSP PGD [36] 1.80 9.71 12.80 48.83 59.57
Res101 SegPGD [17] 0.90 10.64 12.85 60.41 59.48
CosPGD [1] 0.02 14.02 16.41 50.75 61.01
DAG [53] 35.74 23.56 33.92 51.65 61.42
NI [32] 1.65 8.35 10.01 42.51 46.90
DI [54] 2.18 16.94 19.39 50.57 53.31
TI [11] 1.73 25.15 29.84 50.95 57.30
FSPGD (Ours) 2.29 5.96 7.42 36.63 36.91
DV3 PGD [36] 1.74 15.20 16.54 49.92 60.50
Res101 SegPGD [17] 0.63 17.29 18.04 50.18 60.11
CosPGD [1] 0.01 18.83 19.60 50.44 59.91
DAG [53] 36.68 26.70 36.68 52.25 61.29
NI [32] 1.94 14.15 14.91 44.14 48.74
DI [54] 3.99 22.41 23.87 50.91 53.96
TI [11] 2.63 29.58 32.17 51.90 57.19
FSPGD (Ours) 2.03 2.48 3.25 39.82 47.04
@@ C Additional Examples for Qualitative Evaluation
Figure 5: Visualization of clean image, attacked images, and output predictions on Pascal VOC 2012. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 (second row), and PSPNet-Res101 (third row) are used as target models. (a) Clean image, (b) PGD [36], (c) SegPGD [17], (d) CosPGD [1], (e) FSPGD (Ours).
Figure 6: Visualization of clean image, attacked images, and output predictions on Pascal VOC 2012. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 (second row), and PSPNet-Res101 (third row) are used as target models. are used as target models. (a) Clean image, (b) PGD [36], (c) SegPGD [17], (d) CosPGD [1], (e) FSPGD (Ours).
Figure 7: Visualization of clean image, attacked images, and output predictions on Cityscapes. Deeplabv3-Res101 is used as the source model and Deeplabv3-Res50 (second row), Segformer-MiT B0 (third row), and Mask2former-SwinS (fourth row) are used as target models. (a) Clean image, (b) PGD [36], (c) SegPGD [17], (d) CosPGD [1], (e) FSPGD (Ours).
Figure 8: Visualization of clean image, attacked images, and output predictions on Cityscapes. Deeplabv3-Res101 is used as the source model and Deeplabv3-Res50 (second row), Segformer-MiT B0 (third row), and Mask2former-SwinS (fourth row) are used as target models. (a) Clean image, (b) PGD [36], (c) SegPGD [17], (d) CosPGD [1], (e) FSPGD (Ours).
@@ D Detailed Experimental Results for Ablation Studies
This section presents the quantified experimental results in ablation studies discussed in Sec. 4.3 and provides a more detailed explanation of these results. Additionally, it elaborates on ablation study findings that were not included in the main text due to space constraints.
@@ D.1 Performance comparison based on $\tau$ value
The proposed method includes a user-defined parameter, $\tau$, which is used to build the mask $\textrm{M}_{B}$. Since the $\tau$ value affects the attack performance, we conducted extensive experiments to compare the results. As shown in Table 3 and Table 4, the attack performance varies slightly depending on the $\tau$ value. Notably, although performance fluctuates with different $\tau$ values, it consistently outperforms conventional techniques shown in Table 1 and 2 in main paper. We calculated the average performance for each $\tau$ value and selected cos($\pi/3$) as the optimal value, as it achieved the highest average performance.
@@
Table 4: Attack performance comparison in Cityscapes dataset across different $\tau$ values. We measured mIoU scores and bold numbers indicate the best performance for each experimental setup.
[…]
Source Models $\tau$ Source PSP DV3 PSP DV3 Mask2former
Model Res50 Res50 Res101 Res101 Swin-S
Segformer $\pi/6$ 1.70 9.98 16.38 25.98 25.08 43.74
MiT-B0 $\pi/4$ 1.36 11.19 16.59 24.76 24.81 42.97
$\pi/3$ 1.33 10.09 14.57 21.16 22.06 39.92
Source Models $\tau$ Source PSP DV3 PSP DV3 Segformer
Model Res50 Res50 Res101 Res101 MiT-B0
Mask2former $\pi/6$ 4.15 16.83 20.47 26.13 28.58 38.84
Swin-S $\pi/4$ 2.94 16.62 19.61 24.64 26.78 37.80
$\pi/3$ 2.20 15.57 18.00 24.29 25.96 36.87
@@ D.2 Performance comparison based on $\lambda$ value
The proposed loss function consists of two loss terms, $L_{ex}$ and $L_{in}$. Here, we provide a detailed numerical explanation of the experimental results, along with additional results for loss term combinations. Table 3 summarizes the experimental results on the Pascal VOC 2012 and Cityscapes dataset. As shown in Table 5 and 6, the performance of the proposed method varies depending on how the two loss terms are combined. As discussed in the main text, simply adding the two loss terms can result in a compromise, leading to lower performance compared to using $L_{ex}$ alone. To investigate this performance degradation, we conducted experiments with different ratios, such as $L_{ex}+0.5L_{in}$ and $L_{ex}+0.1L_{in}$. The results, as summarized in the Table 5 and 6, show that performance varies depending on the source model; for instance, when PSPNet-Res50 is the source model, performance was lower compared to using $L_{ex}$ alone, but when DeepLabv3-Res50 was used, performance improved. To address this issue of performance variation across source models, we proposed a dynamic $\lambda_{t}$ that adjusts with t, and this method demonstrated the best performance overall.
@@
  Target Models
Source Models $\lambda$ Source Model PSPRes101 DVRes101 FCNVGG16
@@
$\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$ 3.44 21.89 16.57 19.36
Source Models $\lambda$ Source Model PSPRes50 DVRes50 FCNVGG16
PSPRes101 $L_{ex}$ 3.13 14.36 14.68 21.03
$L_{in}$ 4.75 45.67 48.64 37.59
$L_{ex}+L_{in}$ 3.04 12.97 14.57 21.41
$L_{ex}+0.5L_{in}$ 3.06 13.17 14.22 21.08
$L_{ex}+0.1L_{in}$ 3.12 13.83 14.48 21.02
$\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$ 2.99 12.48 13.54 21.30
DV3Res101 $L_{ex}$ 3.32 12.60 13.44 20.57
$L_{in}$ 17.54 65.77 66.57 39.99
$L_{ex}+L_{in}$ 3.71 32.42 30.21 23.37
$L_{ex}+0.5L_{in}$ 3.61 29.84 27.75 22.20
$L_{ex}+0.1L_{in}$ 3.57 27.86 25.17 21.12
$\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$ 3.28 11.42 13.45 21.49
  Target Models
Source Models $\lambda$ Source PSP DV PSP DV Mask2former
Model Res50 Res50 Res101 Res101 Swin-S
Segformer $L_{ex}$ 60.20 55.94 60.08 60.20 64.11 64.62
MiT-B0 $L_{in}$ 21.73 30.56 28.59 36.75 37.91 47.93
$L_{ex}+L_{in}$ 56.43 52.85 56.40 57.59 60.04 63.53
$L_{ex}+0.5L_{in}$ 58.46 54.55 58.78 59.34 62.27 64.17
$L_{ex}+0.1L_{in}$ 59.92 55.61 59.80 60.47 63.78 64.49
$\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$ 1.33 10.09 14.57 21.16 22.06 39.92
Source Models $\lambda$ Source PSP DV PSP DV Segformer
Model Res50 Res50 Res101 Res101 MiT-B0
Maskformer $L_{ex}$ 67.05 58.61 61.84 61.77 64.21 58.58
Swin-S $L_{in}$ 24.05 44.91 44.76 44.77 50.27 51.84
$L_{ex}+L_{in}$ 64.58 55.94 59.34 59.00 62.06 57.42
$L_{ex}+0.5L_{in}$ 65.99 57.40 60.67 60.79 63.28 57.97
$L_{ex}+0.1L_{in}$ 67.43 58.38 61.59 61.87 64.38 58.52
$\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$ 2.20 15.57 18.00 24.29 25.96 36.87
@@ D.3 Performance comparison based on layer location
Unlike conventional methods, the proposed method performs attacks by leveraging intermediate-layer features, making it the first approach to introduce intermediate-layer attacks in the field of semantic segmentation. As such, unlike intermediate-layer attack methods in image classification, there is no prior research on which layer is optimal for attacks in semantic segmentation. To address this, we conducted extensive experiments by attacking various layers of the encoder and summarized the results. Tables 7,  8, and  9 present the intermediate-layer attack performance for ResNet50, ResNet101 encoders, and transformer encoders, respectively. Attacking the later layers of the encoder (i.e., layer 4_2) results in strong performance on the source model but poor performance on the target models. In contrast, attacking the middle layers demonstrates reasonable attack performance on the source model while also achieving high transferability. Therefore, as the proposed method aims to enhance transferability, we chose to attack the middle layers.
@@
Table 9: Attack performance results on source models using the transformer encoder across different attack layers, evaluated on the Cityscapes dataset. We measured mIoU scores and bold numbers indicate the best performance for each experimental setup.
  Target Models
Source Layer name Source PSP DV3 PSP DV3 Mask2former
Models Model Res50 Res50 Res101 Res101 Swin-S
Segformer patch_embeddings-0 43.55 25.45 27.83 30.83 33.77 51.55
MiT-B0 patch_embeddings-1 2.54 10.88 17.94 23.39 25.40 43.25
patch_embeddings-2 1.53 13.56 17.04 27.39 25.47 42.05
patch_embeddings-3 0.67 13.97 15.49 25.53 22.75 44.06
block-0-0 22.11 16.02 23.01 26.31 28.99 51.51
block-0-1 15.28 12.53 18.58 24.94 28.53 46.98
block-1-0 1.33 10.09 14.57 21.16 22.06 39.92
block-1-1 1.16 9.02 13.11 24.03 21.58 43.41
block-2-0 1.01 12.72 13.80 22.98 20.18 40.12
block-2-1 0.86 11.85 13.94 23.82 21.94 41.06
block-3-0 0.33 17.79 18.56 27.58 25.99 45.51
block-3-1 0.16 24.99 25.54 33.59 34.81 50.41
Source Layer name Source PSP DV3 PSP DV3 Segformer
Models Model Res50 Res50 Res101 Res101 MiT-B0
Mask2former embeddings 57.96 27.09 40.40 34.54 44.27 53.06
Swin-S layers-0-blocks-0 48.77 23.98 38.13 29.91 42.71 50.91
layers-0-blocks-1 37.73 16.41 33.17 24.86 36.57 46.63
layers-1-blocks-0 33.11 19.11 25.37 23.72 29.57 44.34
layers-1-blocks-1 17.77 13.65 25.81 19.64 28.03 41.63
layers-2-blocks-0 2.20 15.57 24.29 18.00 25.96 36.87
layers-2-blocks-17 0.79 30.47 36.66 31.20 38.85 45.15
layers-3-blocks-0 1.11 31.94 38.83 34.20 41.06 45.86
layers-3-blocks-1 1.41 33.83 40.14 36.43 43.11 46.49